in 1996 there were aproxmitely 275 students in the Delaware High School band. In 2002, that number increased to 305. Find the annaul rate of change in the number of students in a band.
Answer by stanbon(75887)

in 1996 there were aproxmitely 275 students in the Delaware High School band. In 2002, that number increased to 305. Find the annaul rate of change in the number of students in the band.
--------------
Rate of change = (305-275)/(2002-1996) = 30/6 = 5 students per year
===============
Cheers,
Stan H.